OpenStudy (anonymous):

Two planes are 1,700 miles apart, and their speeds differ by 50 mph. They travel toward each other and meet in 2 hours. Find the speed of the slower plane.

OpenStudy (anonymous):

r t = d Let r be the speed of the first plane. Solve the following for r: {r* 2 + (r - 50)*2 = 1700} r = 450 mph, 450-50 = 400 mph

OpenStudy (anonymous):

a=x b=x+50 2a+2b=1700 2a+2(a+50)=1700 4a+100=1700 a=400

OpenStudy (anonymous):

Thank you thank you!

hero (hero):

Wow. Isn't that a happy face?